The monthly sales for Telco Batteries Inc., were as follows: Month       Sales                        &n

  1. The monthly sales for Telco Batteries Inc., were as follows:

Month       Sales                              

  1.        20
  2.    21
  3.    15
  4.        14
  5.    13
  6.    16
  7.    17
  8.        18
  9.    20
  10.        20
  11.    21
  12.    23

              

  1. Compute a five and a seven-month moving average for months 8 thru 12

  1. Using the data from problem 2:
  1. Determine which of these forecasts are more accurate using the MAD.
  1. How would you explain MAD to your supervisor?
Answer #1


MONTH

SALE

FORMULA

5PMA

FORMULA

7PMA

1

20

2

21

3

15

4

14

5

13

6

16

(20 + 21 + 15 + 14 + 13) / 5 = 16.6

16.6

7

17

(21 + 15 + 14 + 13 + 16) / 5 = 15.8

15.8

8

18

(15 + 14 + 13 + 16 + 17) / 5 = 15

15

(20 + 21 + 15 + 14 + 13 + 16 + 17) / 7 = 16.57

16.57

9

20

(14 + 13 + 16 + 17 + 18) / 5 = 15.6

15.6

(21 + 15 + 14 + 13 + 16 + 17 + 18) / 7 = 16.29

16.29

10

20

(13 + 16 + 17 + 18 + 20) / 5 = 16.8

16.8

(15 + 14 + 13 + 16 + 17 + 18 + 20) / 7 = 16.14

16.14

11

21

(16 + 17 + 18 + 20 + 20) / 5 = 18.2

18.2

(14 + 13 + 16 + 17 + 18 + 20 + 20) / 7 = 16.86

16.86

12

23

(17 + 18 + 20 + 20 + 21) / 5 = 19.2

19.2

(13 + 16 + 17 + 18 + 20 + 20 + 21) / 7 = 17.86

17.86


MAD FOR 5PMA


PERIOD

ACTUAL DEMAND

FORECAST

DEVIATIONS

ABSOLUTE DEVIATION

1

20

2

21

3

15

4

14

5

13

6

16

7

17

8

18

16.57

1.43

9

20

16.29

3.71

3.71

10

20

16.14

3.86

3.86

11

21

16.86

4.14

4.14

12

23

17.86

5.14

5.14

SIGMA

16.85


MAD = SIGMA(ABS DEV) / N

MAD = 16.85 / 5 = 2.84


MAD FOR 7PMA


PERIOD

ACTUAL DEMAND

FORECAST

DEVIATIONS

ABSOLUTE DEVIATION

1

20

2

21

3

15

4

14

5

13

6

16

7

17

8

18

16.57

1.43

9

20

16.29

3.71

3.71

10

20

16.14

3.86

3.86

11

21

16.86

4.14

4.14

12

23

17.86

5.14

5.14

SIGMA

16.85


MAD = 16.85 / 5 = 3.37



2. BASED ON MAD, 5PMA APPEARS TO BE MORE ACCURATE.

3. MAD IS THE REPRESENTATION OF HOW MUCH THE FORECASTED VALUES DEVIATE FROM THE ACTUAL DEMAND, SHOWING ITS DISTRIBUTION BOTH UNDER AND OVER SINCE WE USE ABSOLUTE VALUES TO REPRESENT IT.